OpenStudy (anonymous):

PLEASE HELPP WILL FAN + MEDAL!!! Will medal + fan!!! What is the rate of change? –2x + 4y > 8

OpenStudy (raffle_snaffle):

make it a function of y respect to x

OpenStudy (raffle_snaffle):

x=time and y=the volume or substance

OpenStudy (anonymous):

What? @raffle_snaffle

OpenStudy (anonymous):

Isolate y.

OpenStudy (anonymous):

how do I isolate y @swagmaster47?

OpenStudy (anonymous):

The same way you would do it if it was –2x + 4y = 8

OpenStudy (anonymous):

Just subtract 2x and divide the quantity by 4

OpenStudy (raffle_snaffle):

like you would any function with an equals sign just remember when you are dividing by negative you have to flip the < sign

OpenStudy (anonymous):

You actually don't need to worry about that in this problem, since it would have no effect on the rate of change. @raffle_snaffle

OpenStudy (anonymous):

what do I subtract -2x from?

OpenStudy (anonymous):

both sides

OpenStudy (anonymous):

so I end up with 4y then dividing 4 by 4y isolates y right?

OpenStudy (anonymous):

or wait no i get 2y or no

OpenStudy (raffle_snaffle):

isolate y...

OpenStudy (raffle_snaffle):

you have a question swagmaster?

OpenStudy (anonymous):

-2x + 4y = 8 + 2x +2x 4y = 2x + 8 then divide both sides by 4

OpenStudy (anonymous):

I divide 4y and 2x + 8 by 4 @swagmaster47?

OpenStudy (anonymous):

Yes so y = (1/2)x + 2

OpenStudy (anonymous):

And the rate of change is the slope which is 1/2

OpenStudy (anonymous):

thanks @swagmaster47